# Universal Transient Stability Analysis: A Large Language Model-Enabled Dynamics Prediction Framework

### v1

@@ 4.3 Zero-shot Generalizability on Unseen Faults
This section evaluates the model’s zero-shot generalizability to unseen, severe contingencies (N2N-2 and N3N-3) under mixed stability conditions. As detailed in Table 4 and Table 5, all baseline models exhibited a significant performance collapse when confronted with unseen faults. The $MSE_{H}$ for the RNN-based LSTM and DNR models surged by factors of approximately 13 and 30, respectively, compared to their N1N-1 performance. Even the transformer-based ENC model underperformed, with its $MSE_{U}$ exceeding 1.5. In contrast, TSA-LLM demonstrated superior zero-shot generalization, with its error on unstable samples ($MAE_{U},MSE_{U}$) remaining remarkably stable and increasing by less than 0.01 relative to the N1N-1 scenario. This robustness provides a substantial advantage, with TSA-LLM reducing $MSE_{H}$ by nearly 71.05%–99.17% for N2N-2 settings and 84.27%–99.78% for N3N-3 settings relative to the baselines. These performance disparities are rooted in fundamental architectural and training differences. The failure of DNR, for instance, stems not only from the inherent modeling constraints of its RNN architecture but also from its lack of a mechanism to suppress error propagation, leading to severe error accumulation in complex and unstable scenarios. TSA-LLM incorporates a TeaF and SchS scheme specifically to build resilience to its own predictive errors, ensuring stable long-horizon forecasting. Furthermore, when compared to the ENC baseline, TSA-LLM’s advantage is twofold. Its larger model scale provides greater expressive capacity, while its causal attention mechanism is better aligned with the nature of online TSA tasks than the ENC’s bidirectional attention, explaining its superior performance.
Table 4: New England 39-Bus System: Zero-shot Prediction Capacity Comparison under N-2 Setting
@@
Table 5: New England 39-Bus System: Zero-shot Prediction Capacity Comparison under N-3 Setting
@@ 4.4 Scalability on heterogeneous system
This section evaluates TSA-LLM’s ability to generalize and adapt to a heterogeneous power system with limited target-system data. The model’s inherent cross-system potential was initially evaluated through a direct zero-shot test on the $D_{\text{test}}^{189\text{bus}}$ dataset, as its channel independence design permits direct application to systems of varying dimensions. Despite the significant topological and dynamic differences between the systems, TSA-LLM achieved an impressive $MSE_{H}$ of approximately 0.752. This zero-shot performance is remarkably strong, even surpassing the in-distribution performance of baseline models like LSTM and DNR on their native 39-bus system as shown in Table 3.
Furthermore, to assess the model’s few-shot learning capability, TSA-LLM pre-trained on the 39-bus system was further fine-tuned using varying proportions of $D_{\text{train}}^{189\text{bus}}$, employing the proposed TeaF and SchS schemes. Performance was subsequently evaluated on $D_{\text{test}}^{189\text{bus}}$. As shown in Figure 5, the $MSE_{H}$ of TSA-LLM rapidly deceases with increasing proportions of the 189-bus fine-tuning data and trends for other metrics are generally consistent. For a stringent comparison, the ENC model was trained on the full $D_{\text{train}}^{189\text{bus}}$ dataset to serve as an ”expert” baseline for the target system. The results demonstrate that TSA-LLM achieves performance comparable to this fully-trained expert after fine-tuning on only 4%–5% of the available target-system data. This remarkable data efficiency is visually confirmed by the predicted trajectories from the TSA-LLM fine-tuned by 5% samples from $D_{\text{train}}^{189\text{bus}}$ in Figure 5. Furthermore, when fine-tuned on the complete 189-bus dataset, TSA-LLM reduces the $MSE_{H}$ by nearly 89.62% compared to the expert ENC model, underscoring its superior architectural capacity. This cross-system performance introduces the potential to replace the costly, expert-intensive process of building custom-built models for large-scale systems by enabling an efficient methodology where a foundational model is trained on smaller systems and subsequently adapted with minimal target-system data. Beyond this immediate economic advantage, our findings reveal a clear path for true performance scalability. The model is designed to continually enhance its multi-system, multi-scenario proficiency by incorporating data from an ever-expanding corpus of power systems, paving the way for a single, continually evolving foundation TSA model.
@@
(d) Predicted $\delta$ in stable OC
(e) Predicted $\delta$ in unstable OC
Figure 5: Iceland 189-bus system: few shot scalability. To facilitate presentation, prediction results for a subset of 15 SGs are displayed.
@@ 4.6 Validation on Angular Speed
Recognizing the operational need for comprehensive state variable monitoring, this subsection evaluates the proposed model’s predictive performance for rotor angular speed, with all predictions presented in per-unit (p.u.) values. As summarized in Table 7, the $MSE_{H}$ results indicate that the proposed model demonstrates an even more substantial performance advantage in rotor speed prediction compared to its rotor angle prediction capabilities. Across the four primary evaluation tasks, the proposed model achieves performance improvements of up to two orders of magnitude over baselines. For instance, when trained with the complete $D_{\text{train}}^{\text{189bus}}$, the proposed approach reduced the $MSE_{H}$ by nearly 95.95% relative to the leading baseline model (ENC) on $D_{\text{test}}^{\text{189bus}}$. Furthermore, in a few-shot learning scenario on this system, the proposed method matched the performance of the fully trained ENC model while utilizing only approximately 3% of the training data—a smaller proportion than required for the rotor angle prediction task. These findings underscore the proposed method’s efficacy and versatility in predicting various critical state variables.
Table 7: Angular Speed Prediction Performance Comparison by $MSE_{H}$
@@ 4.7 Training and Inference Cost Comparison
This section analyzes the computational efficiency of the proposed TSA-LLM, comparing it against the ENC baseline and the traditional TDS on the large-scale 189-bus system. We assess efficiency using two key metrics: 1) the training cost per step, defined as a single forward and backward propagation pass on one batch, and 2) the inference time to predict the complete trajectories for a single operating condition (OC). The detailed results of this comparison are presented in Table 8.
@@ 4.7.1 Training Efficiency
The remarkable training efficiency of the large-scale TSA-LLM, which remains comparable to much smaller models, is almost entirely attributable to its parameter-efficient fine-tuning approach. The strategy of freezing the main T-blocks and training only 0.69% of the total parameters significantly reduces the computational complexity associated with the most intensive phase of training, namely the backward pass and parameter update. This allows the model to be fine-tuned with exceptional speed, evidenced by a per-step training time on the large 189-bus system that is only negligibly slower (0.004s) than the compact ENC baseline.
@@ 4.7.2 Inference Speed
In the inference stage, TSA-LLM exhibits a significant speed advantage, proving faster than even the smaller ENC baseline on the 189-bus system. This high efficiency stems from three core design principles. First, the patch and stride operations strategically reduce the input sequence length for the attention mechanism, which is instrumental in mitigating the quadratic computational complexity ($\mathcal{O}(N^{2})$) inherent in long-sequence modeling. Second, the causal attention mechanism results in a sparse attention weight matrix where most elements are zero, thereby simplifying the computational process of aggregating patch representations. Most critically, the online application design reframes the problem by transforming all $n_{x}$ system channels into the batch dimension. This enables the model to process the entire dynamic trajectory for a given operating condition simultaneously while fully leveraging the massive parallel processing capabilities of GPUs on large batches. Consequently, TSA-LLM not only outperforms the ENC model but also achieves an inference speedup of approximately 96.88% over traditional TDS. These results underscore the framework’s strong potential for efficient, real-time trajectory prediction.
@@ 4.8 Generic Validation of Pre-trained Parameters
4.8.1 Validation of Performance Gain from Pre-training
The remarkable universality of TSA-LLM is critically dependent on the knowledge embedded in its pre-trained T-blocks. We verified this through a direct comparison on the $D_{\text{test}}^{N-1}$ dataset between TSA-LLM and a variant whose parameters within T-blocks were randomly sampled from a standard distribution. As presented in Table 9, the model without pre-trained parameters performs poorly, whereas TSA-LLM shows dramatic improvements, including reductions of nearly 86.23% in $MAE_{H}$ and 99.41% in $MSE_{H}$. This contrast underscores that the foundational sequence modeling capabilities of pre-trained T-blocks in GPT, though trained on text, are modality-agnostic and successfully generalize to capture the complex patterns of power system dynamics.
@@ 4.8.2 Explaining Performance Gains
As detailed in Appendix A, pre-training process optimizes T-blocks in GPT model to produce stable and co-directional feature representations. To further validate the transfer of these properties to transient dynamics, two analyses were conducted. First, feature stability was quantified by the average cosine similarity between representations of same patches from the final two transformer layers. The pre-trained model exhibited an average cross-layer similarity of 0.893, significantly higher than the 0.673 from the model without pre-trained parameters, indicating superior representational consistency through deeper layers. Second, feature co-direction was evaluated by the frequency of high cosine similarity (exceeding 0.8) representation pairs within the final layer. As shown in Table 9, the pre-trained model yielded 71.44% high-similarity pairs, substantially exceeding the 46.78% from the model without pre-trained parameters. These combined findings confirm that pre-trained T-blocks retain their generic sequence feature extraction when applied to transient dynamics, empirically explaining the efficacy of the pre-trained T-blocks in TSA tasks.

### v2

@@ 4.3 Zero-shot Generalizability on Unseen Faults
This section evaluates the model’s zero-shot generalizability to unseen, severe contingencies (N2N-2 and N3N-3) under mixed stability conditions. As detailed in Table 5 and Table 6, all baseline models exhibited a significant performance collapse when confronted with unseen faults. The $MSE_{H}$ for the RNN-based LSTM and DNR models surged by factors of approximately 13 and 30, respectively, compablack to their N1N-1 performance. Even the transformer-based ENC model underperformed, with its $MSE_{U}$ exceeding 1.5. In contrast, TSA-LLM demonstrated superior zero-shot generalization, with its error on unstable samples ($MAE_{U},MSE_{U}$) remaining remarkably stable and increasing by less than 0.01 relative to the N1N-1 scenario. This robustness provides a substantial advantage, with TSA-LLM blackucing $MSE_{H}$ by nearly 71.05%–99.17% for N2N-2 settings and 84.27%–99.78% for N3N-3 settings relative to the baselines. These performance disparities are rooted in fundamental architectural and training differences. The failure of DNR, for instance, stems not only from the inherent modeling constraints of its RNN architecture but also from its lack of a mechanism to suppress error propagation, leading to severe error accumulation in complex and unstable scenarios. TSA-LLM incorporates a TeaF and SchS scheme specifically to build resilience to its own pblackictive errors, ensuring stable long-horizon forecasting. Furthermore, when compablack to the ENC baseline, TSA-LLM’s advantage is twofold. Its larger model scale provides greater expressive capacity, while its causal attention mechanism is better aligned with the nature of online TSA tasks than the ENC’s bidirectional attention, explaining its superior performance.
Table 5: New England 39-Bus System: Zero-shot Pblackiction Capacity Comparison under N-2 Setting
@@
Table 6: New England 39-Bus System: Zero-shot Pblackiction Capacity Comparison under N-3 Setting
@@ 4.4 Scalability on heterogeneous system
This section evaluates TSA-LLM’s ability to generalize and adapt to a heterogeneous power system with limited target-system data. The model’s inherent cross-system potential was initially evaluated through a direct zero-shot test on the $D_{\text{test}}^{189\text{bus}}$ dataset, as its channel independence design permits direct application to systems of varying dimensions. Despite the significant topological and dynamic differences between the systems, TSA-LLM achieved an impressive $MSE_{H}$ of approximately 0.752. This zero-shot performance is remarkably strong, even surpassing the in-distribution performance of baseline models like LSTM and DNR on their native New England 39-bus system, as shown in Table 4.
Furthermore, to assess the model’s few-shot learning capability, TSA-LLM pre-trained on the New England 39-bus system was further fine-tuned using varying proportions of $D_{\text{train}}^{189\text{bus}}$, employing the proposed TeaF and SchS schemes. Performance was subsequently evaluated on $D_{\text{test}}^{189\text{bus}}$. As shown in Fig. 6, the $MSE_{H}$ of TSA-LLM rapidly decreases with increasing proportions of the Iceland 189-bus system fine-tuning data, and trends for other metrics are generally consistent. For a stringent comparison, the ENC model was trained on the full $D_{\text{train}}^{189\text{bus}}$ dataset to serve as an ”expert” baseline for the target system. The results demonstrate that TSA-LLM achieves performance comparable to this fully-trained expert after fine-tuning on only 4%–5% of the available target-system data. This remarkable data efficiency is visually confirmed by the pblackicted trajectories from the TSA-LLM fine-tuned by 5% samples from $D_{\text{train}}^{189\text{bus}}$ in Fig. 6 and in Fig. 7, where pblackictions and ground truths for SG 2 are plotted on the same axes for improved clarity. Furthermore, when fine-tuned on the complete Iceland 189-bus system dataset, TSA-LLM blackuces the $MSE_{H}$ by nearly 89.62% compablack to the expert ENC model, underscoring its superior architectural capacity. This cross-system performance introduces the potential to replace the costly, expert-intensive process of building custom-built models for large-scale systems by enabling an efficient methodology where a foundational model is trained on smaller systems and subsequently adapted with minimal target-system data. Beyond this immediate economic advantage, our findings reveal a clear path for true performance scalability. The model is designed to continually enhance its multi-system, multi-scenario proficiency by incorporating data from an ever-expanding corpus of power systems, paving the way for a single, continually evolving foundation TSA model.
@@
(d) Pblackicted $\delta$ in stable OC
(e) Pblackicted $\delta$ in unstable OC
Fig. 6: Iceland 189-bus system: few shot scalability. To facilitate presentation, pblackiction results for a subset of 15 SGs are displayed.
(a) Stable OC
(b) Unstable OC
Fig. 7: Representative $\delta$ trajectory comparison for SG 2 (189-bus system). Pblackictions and ground truths are plotted on the same axes for improved clarity.
@@ 4.6 Validation on Angular Speed
Recognizing the operational need for comprehensive state variable monitoring, this subsection evaluates the proposed model’s pblackictive performance for rotor angular speed, with all pblackictions presented in per-unit (p.u.) values. As summarized in Table 8, the $MSE_{H}$ results indicate that the proposed model demonstrates an even more substantial performance advantage in rotor speed pblackiction compablack to its rotor angle pblackiction capabilities. Across the four primary evaluation tasks, the proposed model achieves performance improvements of up to two orders of magnitude over baselines. For instance, when trained with the complete $D_{\text{train}}^{\text{189bus}}$, the proposed approach blackuced the $MSE_{H}$ by nearly 95.95% relative to the leading baseline model (ENC) on $D_{\text{test}}^{\text{189bus}}$. Furthermore, in a few-shot learning scenario on this system, the proposed method matched the performance of the fully trained ENC model while utilizing only approximately 3% of the training data—a smaller proportion than requiblack for the rotor angle pblackiction task. These findings underscore the proposed method’s efficacy and versatility in pblackicting various critical state variables.
To further explore alternative data processing strategies, an additional experiment, TSA-LLM (Unit-wise), is conducted, as shown in the last row of Table 8. While the original TSA-LLM processes state variables independently to maximize cross-system adaptability, we investigate an alternative approach inspiblack by [34], where each synchronous generator (SG) is treated as an integrated unit. In this unit-wise formulation, the state trajectories of each SG (e.g., rotor angle $\delta$ and speed $\omega$) are concatenated and jointly represented as multivariate time series, enabling the model to capture intra-unit correlations while maintaining inter-system generality. As shown in Table 8, the unit-wise variant achieves a $95.33\%$ blackuction in $\mathrm{MSE}_{H}$ on the full $D_{\text{test}}^{189\text{bus}}$ and matches the fully trained ENC baseline using only $3\%$ of the training data. These results demonstrate that unit-wise temporal modeling provides a viable alternative for capturing variable dependencies while preserving cross-system flexibility.
Table 8: Angular Speed Pblackiction Performance Comparison by $MSE_{H}$
@@ 4.7 Training and Inference Cost Comparison
This section analyzes the computational efficiency of the proposed TSA-LLM, comparing it against the ENC baseline and the traditional TDS on the large-scale 189-bus system. We assess efficiency using two key metrics: 1) the training cost per step, defined as a single forward and backward propagation pass on one batch, and 2) the inference time to pblackict the complete trajectories for a single operating condition (OC). The detailed results of this comparison are presented in Table 9.
@@ 4.7.1 Training Efficiency
The remarkable training efficiency of the large-scale TSA-LLM, which remains comparable to much smaller models, is almost entirely attributable to its parameter-efficient fine-tuning approach. The strategy of freezing the main T-blocks and training only 0.69% of the total parameters significantly blackuces the computational complexity associated with the most intensive phase of training, namely the backward pass and parameter update. This allows the model to be fine-tuned with exceptional speed, evidenced by a per-step training time on the large 189-bus system that is only negligibly slower (0.004s) than the compact ENC baseline.
@@ 4.7.2 Inference Speed
In the inference stage, TSA-LLM exhibits a significant speed advantage, proving faster than even the smaller ENC baseline on the 189-bus system. This high efficiency stems from three core design principles. First, the patch and stride operations strategically blackuce the input sequence length for the attention mechanism, which is instrumental in mitigating the quadratic computational complexity ($\mathcal{O}(N^{2})$) inherent in long-sequence modeling. Second, the causal attention mechanism results in a sparse attention weight matrix where most elements are zero, thereby simplifying the computational process of aggregating patch representations. Most critically, the online application design reframes the problem by transforming all $n_{x}$ system channels into the batch dimension. This enables the model to process the entire dynamic trajectory for a given operating condition simultaneously while fully leveraging the massive parallel processing capabilities of GPUs on large batches. Consequently, TSA-LLM not only outperforms the ENC model but also achieves an inference speedup of approximately 96.88% over traditional TDS. These results underscore the framework’s strong potential for efficient, real-time trajectory pblackiction.
It is worth clarifying that the reported 0.57 s corresponds to the end-to-end inference time under our specific hardware and implementation settings for pblackicting the complete post-fault trajectory over a 9 s horizon for one OC, and it is not a fundamental lower bound. In practical deployment, standard GPU inference optimizations (e.g., mixed precision/quantization and optimized runtime backends) can further blackuce latency. Moreover, for many unstable cases, system instability can often be identified from the early-stage response without requiring full-horizon pblackiction, implying that the effective response time for triggering emergency control can be substantially shorter than 0.57 s when a shorter screening horizon is adopted. Finally, our online application design batches all channels for concurrent pblackiction via GPU parallelization, so the inference time does not scale linearly with system size in the same way as time-domain simulation; as system scale increases and TDS becomes slower, the relative speed advantage of TSA-LLM is expected to be even more pronounced.
@@ 4.8 Generic Validation of Pre-trained Parameters
4.8.1 Performance Gain Attributable to Pre-training
Pre-trained T-blocks are a key factor behind the cross-scenario generalization of TSA-LLM. To quantify their contribution, we compablack TSA-LLM with a counterpart in which the T-block parameters were randomly initialized, using the same architecture and evaluation setting on $D_{\text{test}}^{N-1}$. As shown in Table 10, the randomly initialized model exhibits substantial performance degradation, while TSA-LLM achieves markedly lower errors, with about $86.23\%$ and $99.41\%$ blackuctions in $MAE_{H}$ and $MSE_{H}$, respectively. These results indicate that pre-trained T-blocks within LLMs provide transferable sequence modeling priors for capturing transient power-system dynamics.
@@ 4.8.2 Explaining Performance Gains
After establishing the performance gap, we further investigated whether pre-trained T-blocks preserve the representational properties described in A. To align this analysis with the subsequent mechanism discussion, we used two indicators: feature stability and feature co-directionality (see Table 10).
Feature stability characterizes cross-layer consistency, i.e., whether the same patch remains similarly encoded as depth increases. We measure it by the mean cosine similarity between the representations of each patch in the last two Transformer layers. TSA-LLM attains 0.893, compablack with 0.673 for the randomly initialized counterpart, indicating more stable feature propagation across layers.
Feature co-directionality characterizes intra-sequence alignment at a fixed layer. For each sequence, we compute pairwise cosine similarities among patch representations in the final layer and report the proportion of patch pairs with cosine similarity larger than 0.8. TSA-LLM reaches 71.44%, versus 46.78% without pre-training.
Overall, these results support the mechanism analysis in A and indicate that pre-trained T-blocks preserve transferable sequence representation properties for transient dynamics modeling. From a physical perspective, transient trajectories in power systems with synchronous generators are governed by the same basic physics mechanisms, e.g., the swing equation coupled with network algebraic constraints. Therefore, after appropriate normalization, different systems often exhibit similar temporal evolution patterns despite differences in network scale and parameters. The stable temporal representations extracted by pre-trained T-blocks are effective in capturing these shablack patterns, which helps explain their strong performance in cross-scenario and cross-system generalization.
@@
4.9 Backbone Replacement with Qwen3-0.6B
To further validate that the proposed framework is backbone agnostic, we instantiated TSA-LLM with an alternative pretrained decoder-only Transformer, Qwen3-0.6B. While both the original GPT backbone and Qwen3-0.6B adopt a decoder-only architecture, Qwen3-0.6B additionally uses a Mixture-of-Experts design. For a controlled comparison, we kept the data processing pipeline unchanged, including channel-independence, sample-wise normalization, and patching. We also used the same training protocol, including TeaF and SchS, and the same evaluation setup. In other words, we changed the backbone while holding all other components fixed.
Table 11 reports the resulting $MSE_{H}$ scores. The Qwen3-0.6B (MoE) variant yields consistently lower errors than the original TSA-LLM across all four benchmarks, including $D_{\text{test}}^{N-1}$, $D_{\text{test}}^{N-2}$, $D_{\text{test}}^{N-3}$, and $D_{\text{test}}^{\text{189bus}}(5\%)$.
These results lead to two observations. First, the TSA-LLM framework transfers naturally to other pretrained large-model backbones beyond GPT within the decoder-only family, which supports its architectural compatibility and portability. Second, increasing the effective model capacity via an MoE backbone appears to strengthen sequence representations and nonlinear dynamics modeling, improving long-horizon trajectory pblackiction in both unseen-fault and cross-system settings. Overall, scaling the backbone provides a practical route to further enhance TSA universality and pblackiction accuracy.
Table 11: Estimated $MSE_{H}$ comparison after backbone replacement with Qwen3-0.6B
Model $D_{\text{test}}^{N-1}$ $D_{\text{test}}^{N-2}$ $D_{\text{test}}^{N-3}$ $D_{\text{test}}^{\text{189bus}}(5\%)$
TSA-LLM (GPT) 0.137 0.163 0.154 0.156
TSA-LLM (Qwen3-0.6B MoE) 0.126 0.149 0.141 0.142
4.10 Cross-System Scaling on the IEEE 68-bus System and the IEEE 118-bus System
To further strengthen the empirical justification of universality within the IEEE benchmark regime and to better illustrate the scaling trend beyond the IEEE-39/189 setting, we additionally evaluate TSA-LLM on two additional benchmark systems, namely the IEEE 68-bus system and the IEEE 118-bus system, under N1N\!-\!1 contingencies. The transient trajectories are generated following the same simulation pipeline described in Sec. 4.1 (Data Generation and Model Settings). This experiment is intended to provide additional benchmark evidence on cross-system scalability and data-efficiency, rather than to claim that IEEE-39 data alone can guarantee direct zero-shot generalization to real utility-scale grids with substantially higher heterogeneity and renewable penetration. Specifically, we generate 14,400 trajectories for the IEEE 68-bus system (16 SGs) and 17,100 trajectories for the IEEE 118-bus system (19 SGs), corresponding to 900 operating conditions (OCs) in total, and split each dataset into $D_{\text{train}}^{\text{target}}$ and $D_{\text{test}}^{\text{target}}$ with a 2:1 ratio (9,600/4,800 for the IEEE 68-bus system and 11,400/5,700 for the IEEE 118-bus system). In particular, the test set comprises 300 OCs, and the OCs are generated with a stable/unstable proportion of 7:3.
In addition to trajectory-pblackiction errors, we further report stability-relevant classification metrics derived from the pblackicted trajectories. Specifically, we define a transient stability index (TSI) based on the maximum rotor-angle separation with respect to the center-of-inertia (COI) [29, 28]. For an OC $k$, let $\delta_{i,k}(t)$ denote the rotor angle of generator $i$, and define
$\Delta\delta_{k}(t)$ $=\max_{i}\left|\delta_{i,k}(t)-\delta_{\text{COI},k}(t)\right|,$ (26)
$\mathrm{TSI}_{k}$ $=\max_{t\in[t_{0},\,T]}\Delta\delta_{k}(t),$ (27)
where $t_{0},T]$ is the post-fault simulation horizon. Given a pblackefined stability threshold $\delta_{\text{th}}$ (set to $\pi$), the ground-truth stability label is
$y_{k}=\mathbb{I}\left(\mathrm{TSI}_{k}>\delta_{\text{th}}\right)\in\{0,1\},$ (28)
where $y_{k}=1$ indicates an unstable (loss-of-synchronism) OC and $y_{k}=0$ indicates a stable OC. The pblackicted label $\hat{y}_{k}$ is obtained by applying the same rule to the pblackicted trajectories. We then report the classification accuracy (Acc) and the recall of the unstable class (Recall-U) as
$\mathrm{Acc}$ $=\frac{1}{N}\sum_{k=1}^{N}\mathbb{I}\left(\hat{y}_{k}=y_{k}\right),$ (29)
$\mathrm{Recall\text{-}U}$ $=\frac{\sum_{k=1}^{N}\mathbb{I}\left(\hat{y}_{k}=1\right)\mathbb{I}\left(y_{k}=1\right)}{\sum_{k=1}^{N}\mathbb{I}\left(y_{k}=1\right)},$ (30)
where $N$ is the number of OCs in the test set.
We consider three transfer settings: (i) zero-shot evaluation, where the model is fine-tuned only on the IEEE 39-bus system and directly tested on the target system; (ii) few-shot adaptation on the target system with limited data ratios (10% and 20%); and (iii) full-data fine-tuning on the target system. For comparison, we also report the performance of the strongest baseline, ENC, trained with the full target-system data as an expert model.
Table 12 summarizes the results. On the IEEE 68-bus system, the zero-shot transfer from the IEEE 39-bus system attains an $MSE_{H}$ of 0.620, and few-shot fine-tuning with only 10% target data already outperforms the full-data ENC expert (0.122 vs. 0.135). With 20% target data, the error is further blackuced to 0.106, and full-data fine-tuning reaches 0.082. A similar trend is observed on the IEEE 118-bus system, where $MSE_{H}$ decreases from 0.710 (zero-shot) to 0.136 with 10% target data, surpassing the ENC expert trained on the full target data (0.149), and further improves to 0.113 (20%) and 0.089 (full-data). Notably, the conventional RNN baselines (LSTM and DNR) exhibit orders-of-magnitude larger trajectory errors and markedly lower unstable-recall, indicating their limited robustness in mixed stable/unstable post-fault settings.
For a supplementary comparison on the shablack task of transient stability classification, we further include the Potential Energy Boundary Surface (PEBS) method as a representative classical direct-method baseline [7]. PEBS is an energy-function-based method that approximates the post-fault stability boundary by locating the first local maximum of the potential energy along the fault-on trajectory, thereby avoiding explicit UEP computation and retaining attractive screening efficiency for large contingency sets. Since PEBS outputs only stability decisions rather than pblackicted trajectories, $MSE_{H}$ is not applicable. As shown in Table 12, PEBS attains Acc/Recall-U of 75.0%/76.7% on the IEEE 68-bus system and 77.7%/67.8% on the IEEE 118-bus system. Its unstable-class recall remains higher than those of the conventional RNN baselines, showing that PEBS can still capture a meaningful portion of critical unstable cases. However, its relatively modest overall accuracy suggests a conservative decision tendency, i.e., improved unstable-case coverage is achieved at the cost of more false alarms on stable OCs. By contrast, TSA-LLM consistently delivers both higher Acc and higher Recall-U, indicating a more balanced and reliable stability discrimination capability on these benchmark systems, while simultaneously providing full trajectory-level pblackiction and cross-system transfer capability.
Overall, these benchmark results indicate that TSA-LLM can be efficiently adapted to other benchmark systems with limited target-system data, while maintaining a clear improvement trend as the fine-tuning data increases. Extending the same framework to real utility-scale systems will require expanding the training corpus to cover larger-scale and renewable-rich dynamics, which we identify as future work.
Table 12: Cross-system scaling evaluation on the IEEE 68-bus system and the IEEE 118-bus system under N1N\!-\!1 contingencies.
Target system Method/Setting $MSE_{H}$ Acc (%) Recall-U (%)
IEEE 68-bus TSA-LLM (Zero-shot) 0.620 80.0 70.0
TSA-LLM (Few-shot, 10%) 0.122 95.0 95.6
TSA-LLM (Few-shot, 20%) 0.106 96.0 96.7
TSA-LLM (Full-data) 0.082 97.0 97.8
ENC (Full-data) 0.135 94.7 91.1
LSTM (Full-data) 12.876 89.0 55.6
DNR (Full-data) 9.516 90.3 60.0
PEBS - 75.0 76.7
IEEE 118-bus TSA-LLM (Zero-shot) 0.710 78.0 68.9
TSA-LLM (Few-shot, 10%) 0.136 94.3 94.4
TSA-LLM (Few-shot, 20%) 0.113 95.3 96.7
TSA-LLM (Full-data) 0.089 96.3 97.8
ENC (Full-data) 0.149 94.0 90.0
LSTM (Full-data) 15.083 88.0 51.1
DNR (Full-data) 11.809 89.3 55.6
PEBS - 77.7 67.8